# ProGeny I: a new simple stellar population generator and impact of isochrones / stellar atmospheres / initial mass functions

### v2

@@ 4.1.1 Internal Comparisons
A useful top-level comparison between different SSPs is to analyse how much mass is required to produce a given bolometric luminosity, i.e. $M/L$ where both quantities are computed in solar units. Figure 7 compares to $M/L$ for various combinations of ProGeny isochrones and IMFs (using the C3K stellar atmosphere and solar metallicity), with BC03 (using a Chabrier IMF) added for reference. The bottom panels divides through by the BC03 SSP $M/L$ to make variations clearer. The main note is that most combinations of ProGeny isochrones and IMFs produce more mass per unit light, the exception being PARSEC with a Chabrier IMF. In principle that combination should be the most similar to BC03, but we observe that PARSEC is a significant evolution of the older Padova isochrones it was developed from (as used by BC03) and there are other treatments in the BC03 SSP that handle various stellar population aspects differently: remnant stars; planetary nebulae; AGB phases; stellar atmospheres (especially for hot stars) etc.
@@
Figure 7: Comparison of the bolometric mass-to-light ration for different combinations of isochrones and IMF (there is very little variation due to atmosphere choice, so that is left as C3K for this comparison). Top panel shows the measured $M/L$, and bottom panel shows this relative to BC03, where in general ProGeny SSPs produce more mass-to-light than BC03.
A core difference between SSPs is how much their spectra differ at a given epoch. One way to capture this information is to compute the relative integrated absolute flux difference between two spectra as a function of time. This has the desired effect of treating the spectra symmetrically, both penalising the relative under and over production of flux at all wavelengths in a consistent manner. For ProGeny this is presented in Figure 8, where we compare SSPs generated for the MIST, PARSEC, and BaSTI isochrones; C3K, PHOENIX (Husser), MILES, and BaSeL stellar atmospheres; and also the Chabrier, Kroupa, and Salpeter IMFs commonly used in modern literature.
Figure 8: SED comparisons of different isochrones, stellar atmospheres and IMF choices with ProGeny. The y-axis is the relative total integrated flux discrepancy when switching isochrones, stellar atmospheres or IMFs as a function of stellar age, where both relative excesses of flux and deficits are equally penalised. In all cases the major properties kept constant during comparisons are shown in the top left.
The immediately notable feature of Figure 8 is that our choice of isochrones has a much more dramatic impact than our choice of stellar atmospheres or IMF at nearly all stellar ages, with particularly dramatic differences seen at periods of stellar evolution that are both extremely luminous and difficult to model. At around 1 Gyr the big spike is dominated by differences in how the various AGB phases are treated. At much older ages we see the relative flux disagreement settles down to around 12-13%. Since most stellar mass measurements are dominated by stellar population in this age window it suggests fitting error of no less than $\sim 0.05$ dex even in a situation where every other aspect of the process (stellar atmospheres, interpolation, IMF, SED engine) is identical. In comparison, the impact of the IMF when switching between the two most popular modern variants (Chabrier and Kroupa) is extremely minimal, causing only 6% relative flux disagreement and 0.03 dex stellar mass variance. The switch to Salpeter does cause a dramatic spectral shift, but this is because it is extremely bottom-heavy (a large fraction of mass in low mass stars), meaning it produces much less light for a given unit of star formation mass.
@@
Assuming we using either a Chabrier or Kroupa IMF, extremely young ($10^{7}$ Gyr) and old ($10^{10}$ Gyr) stellar population spectra uncertainties are mildly dominated by stellar atmosphere choice (but isochrones are not far behind). The scatter between these is minimised at a few hundred Myr, and slowly rises beyond that to a typical 5–7 % variation. The least disagreement is seen between MILES and BaSeL stellar atmospheres, likely because both of these are empirically constructed. C3K agrees more with these two empirical libraries than PHOENIX (Husser). Given the excellent parameter coverage of C3K (intrinsically much larger and well sampled than MILES or BaSeL), this result partly informs our decision to recommend C3K as our fiducial ‘base’ stellar atmosphere library. As with the isochrones above, SSPs using all of the discussed stellar atmospheres have been pre-generated with ProGeny. Anybody carrying out a detailed study of SEDs with ProSpect should consider using the other stellar atmosphere combinations too.
An obvious issue from an inference perspective is we do not know the truth re isochrones, stellar atmospheres, and/or IMF, and it is quite possible that a combination of the ‘wrong’ isochrone and ‘wrong’ IMF combine in a manner that provides a better overall fit to our data. The above findings suggest that much better fidelity in isochrones and atmospheric modelling would be needed before strong statements about IMF behaviour could be conclusively demonstrated in any absolute sense. Within a consistent SED software environment (e.g. ProSpect) relative statements about the impact of the IMF can be made however. The impact of using different SSPs (including our new ProGeny ones) is explored in depth in our Bellstedt & Robotham (in prep.) companion paper.
@@
Figure 9: Stellar mass remaining comparison of different isochrone and IMF choices with ProGeny(using solar metallicity). Other versions of this Figure (for different metallicities) can be viewed at rpubs.com/asgr.
@@
Figure 10: Ratio of stellar mass remaining for different isochrone and IMF choices (using solar metallicity).
4.1.2 Literature Comparisons
@@
Table 4: Basic information regarding the literature SSPs we compare against. Note for the stellar atmospheres only the dominant library is listed, nearly all SSPs use a complex mixture of spectra to achieve adequate temperature and effective surface gravity parameter coverage. Note M05 only has a Kroupa IMF variant.
@@
BPASS STARS BaSeL Chabrier 6 – 11 51 -3.3 – 0.3 13 Stanway & Eldridge (2018)
@@
The most fundamental comparison we can make between SSPs is by analysing the actual spectra produced for specific combinations of metallicity and stellar evolutionary age. Figue 11 presents such spectra for solar metallicities, covering 100,000 yr through to 10 Gyr. For reasons not entirely clear, FSPS when run with MIST isochrones produces very unusual spectra for ages below $10^{7}$ years, with a much flatter spectral shape than seen for any other SSP. These unusual features are not present when running in the same manner but specifying Padova isochrones with FSPS (using the original Fortran software), so it would seem to be a specific issue with the younger MIST isochrones. For that reason we prefer to present FSPS using Padova isochrones for the following comparisons.
Figure 11: Comparison of solar metallicity spectra at different stellar population ages for different SSPs. EMILES is only shown for ages $10^{7}$ years and older (it does not produce SSPs younger than $\sim$30m years. The ages presented are log spaced with 1 dex intervals, covering 100,000 Years through to 10 Gyr. Other versions (for different metallicities) of this Figure can be viewed at rpubs.com/asgr.
@@
A consistent observation is that predictions for SSP spectra below 1,000 Å are highly variable between specific SSPs for age much above a million years, with sometimes many dex of disagreement re the detail of spectra in this regime. The reason for this is complex, but the general issue is the differing treatment of hot remnant stars, horizontal branch (HB) stars and helium burning phases by different isochrones and/or SPLs. This is clear when observing the M05 SSP that has no remnant stars and coarse treatment of HB stars: we see a sharp truncation of UV light for all SSPs $10^{8}$ years and older with no evidence of any UV upturn at all. It is relatively rare that these parts of the SED are interrogated with lower redshift surveys since even UV facilities like GALEX only probe down to 1350 Å in the FUV. However, by redshift 0.3 the rest frame GALEX FUV filter starts to cover the 1,000 Å feature seen here, and much beyond this even typical optical filters will be measuring flux in this region.
@@
Redder than this 1000 Å feature the SSPs are remarkably consistent over 5 dex in age, with the M05 SSPs being the most obvious outlier below the 4000 Å break. Beyond that observation, the next most significant variation is seen in the near infrared (1 micron and redder) where the various AGB phases become important for ages older than $\sim 10^{7}$ years (especially above $10^{8}$ years) and have complex modelling differences at the isochrones level (as noted in Section 3.1) and decisions regarding the best (or perhaps more accurately, least bad) method for matching these stellar evolution phases to the most appropriate stellar atmospheres. Given the huge wealth of data, it is not surprising that models collectively agree best in the optical regime for giga-year and older stellar populations. This represents the dominant spectral feature in a huge number of passive galaxies and globular clusters, and is a common target when checking the validity of a given SSP.
@@
An additional difference we note between ProGeny (with MIST isochrones) and other SSPs can be seen in the lowest age (100,000 years) top-left panel. Red-wards of about 1 micron there is a significant (almost 1 dex) excess to the classic black-body power-law slope you might expect (and as is seen for other SSPs). This feature is entirely cause by the tracking of pre-main sequence stars in MIST, which produces an additional cool ($\sim 2,000$ K) spectrum representing the thermal emission from collapsing molecular clouds and proto-stars. This cool spectrum is added onto the dominate spectrum of extremely hot O and B stars on the main sequence which have typical temperatures $\sim 50,000$ K. While it looks dramatic when plotted like this, the actual flux contribution in a typical galaxy SED is very small, and this difference in behaviour has little impact on composite (or complex) stellar population spectra in general.
@@
Figure 12: Stellar mass remaining comparison of different isochrone and IMF choices with ProGenyversus a compendium of literature SSPs (using solar metallicity).
@@ 5 Conclusions And Future Outlook
Here we have, in brief, introduced the new R based ProGeny SPL package that allows for the rapid and flexible generation of SSPs for use in particular with ProSpect, but potentially any other SED generating or fitting software. The package is thoroughly documented, with relevant examples for all user-level functions, and longer form vignettes demonstrating sophisticated types of analysis and comparison101010rpubs.com/asgr.
@@
Basic comparisons show that ProGeny is spectrally consistent with other popular SPLs when using similar isochrones and stellar atmospheres, with understandable variations in a few key domains (e.g. UV upturn) due to our ProGeny isochrones incorporating different physical processes to some published SPLs, even nominally identical ones. This is not surprising since isochrones with the same name (e.g. MIST, PARSEC, BaSTI etc) receive regular updates to their physics over time, and the literature is lacking in clear version control for SPLs and/or SSPs. The largest variations in spectral output are noted when switching isochrones ($\sim 60$%), followed by the stellar atmospheres ($\sim 20$%) and finally IMFs ($\sim 10$%, between Chabrier and Kroupa variants).
@@
To create a fiducial set of stellar atmospheres using the recommended combination of MIST for the isochrones, C3K (for the base stellar atmospheres) and a Chabrier IMF takes 5s at low resolution (matching BC03) and 18s at high resolution (matching CB19) using a modern Apple MacBook. These speeds make it realistic to create SSPs on-the-fly, and for many use cases (since the fitting time will dominate the processing time with ProSpect) there is no need to pre-generate static SSPs. This is particularly useful when exploring variations of the input ingredients (like the IMF) since exhaustively generating all potential combinations is onerous in terms of disk space and book-keeping.
@@
As part of this work, we introduce a new standardised FITS based format for defining all the characteristics of stellar atmospheres to be incorporated into a new SPL/SSP. We noted while constructing the default input stellar atmospheres that every single library used its own format, ranging from pure binaries, ASCII text files, and FITS files of many different formats. While it is true we have added ‘yet another standard’, the one we propose is a minimal combination of the best characteristics of the data we processed in a highly open and well used binary (FITS) format.
@@
A companion paper (Bellstedt & Robotham in prep.) will investigate the impact of varying the available ProGeny isochrones, stellar atmospheres, and IMFs when using the SSPs with ProSpect to fit GAMA galaxies. Longer term the aim is to establish the most plausible fiducial model, i.e. what combination of isochrones, stellar atmospheres and IMF (including evolving varieties) best reproduce a self consistent model of galaxy evolution in the Universe. Key to this will be investigating the implied cosmic star formation history and metallicity evolution implied by fitting galaxies at low (GAMA) and moderate (WAVES) redshift, and comparing this to ‘in-situ’ measures at earlier epochs (heavily inspired by Bellstedt et al., 2020).
@@
Planned enhancements for ProGeny include the separate treatment of remnant stars in order to make better use of isochrones that do not track these phases in detail. Until these additional tracks have been incorporated, general purpose SED modelling should probably use the MIST isochrones only. We also plan to incorporate a wider variety of isochrones libraries, where the impact of stellar rotation, $\alpha$ enhancement and potentially binary evolution (with BPASS; Stanway & Eldridge, 2018) can be explored in systematic detail. To self-consistently account for more exotic forms of $\alpha$ enhancement the appropriate stellar atmospheres would also need to be added, which means such extensions are likely to happen slowly (multiple years). The ultimate hope is many of these extensions will come from the wider community, with all aspects of ProGeny being available under a permissive L-GPL3 license.
Data and Software Availability
@@
  • Long-form examples (vignettes) that generate many of the Figures presented in this work available at ASGR RPubs111111rpubs.com/asgr.
  • Script to generate all 45 SSPs used in this work, also at ASGR RPubs.
  • Seven pre-generated ProSpect format (FITS) SSPs that use MIST isochrones, C3K base stellar atmospheres, and the following IMFs:
@@
    • Metallicity Evolving Kroupa (0.1 – 100 M; default ProGeny evolution arguments)
@@
    • Metallicity Evolving Lacey (0.1 – 100 M; default ProGeny arguments)

### v3

@@ 4.2 Internal High-Level Comparisons
A useful high-level comparison between different SSPs is to analyse how much mass is required to produce a given bolometric luminosity, i.e. $M/L$ where both quantities are computed in solar units. Figure 10 compares to $M/L$ for various combinations of ProGeny isochrones and IMFs (using the C3K stellar spectrum and Z), with BC03 (using a Chabrier IMF) added for reference. The bottom panels divides through by the BC03 SSP $M/L$ to make variations clearer. The main note is that most combinations of ProGeny isochrones and IMFs produce more mass per unit light, the exception being PARSEC with a Chabrier IMF. In principle that combination should be the most similar to BC03, but we observe that PARSEC is a significant evolution of the older Padova isochrones it was developed from (as used by BC03) and there are other treatments in the BC03 SSP that handle various stellar population aspects differently: remnant stars; planetary nebulae (explicitly added in BC03, and not currently generated in ProGeny); AGB phases; stellar spectra (especially for hot stars) etc.
@@
Figure 10: Comparison of the bolometric mass-to-light ration for different combinations of isochrones and IMF (there is very little variation due to stellar spectrum choice, so that is left as C3K for this comparison). Top panel shows the measured $M/L$, and bottom panel shows this relative to BC03, where in general ProGeny SSPs produce more mass-to-light than BC03.
A core difference between SSPs is how much their spectra differ at a given epoch. One way to capture this information is to compute the relative integrated absolute flux difference between two spectra as a function of time. This has the desired effect of treating the spectra symmetrically, both penalising the relative under and over production of flux at all wavelengths in a consistent manner. For ProGeny this is presented in Figure 11, where we compare SSPs generated for the MIST, PARSEC, and BaSTI isochrones; C3K, PHOENIX (Husser), MILES, and BaSeL stellar spectra; and also the Chabrier, Kroupa, and Salpeter IMFs commonly used in modern literature.
Figure 11: SED comparisons of different isochrones, stellar spectra and IMF choices with ProGeny. The y-axis is the relative total integrated flux discrepancy when switching isochrones, stellar spectra or IMFs as a function of stellar age, where both relative excesses of flux and deficits are equally penalised. In all cases the major properties kept constant during comparisons are shown in the top left.
The immediately notable feature of Figure 11 is that our choice of isochrones has a much more dramatic impact than our choice of stellar spectra or IMF at nearly all stellar ages, with particularly dramatic differences seen at periods of stellar evolution that are both extremely luminous and difficult to model. At around 1 Gyr the big spike is dominated by differences in how the various AGB phases are treated. At much older ages we see the relative flux disagreement settles down to around 12-13%. Since most stellar mass measurements are dominated by stellar population in this age window it suggests fitting error of no less than $\sim 0.05$ dex even in a situation where every other aspect of the process (stellar spectra, interpolation, IMF, SED engine) is identical. In comparison, the impact of the IMF when switching between the two most popular modern variants (Chabrier and Kroupa) is extremely minimal, causing only 6% relative flux disagreement and 0.03 dex stellar mass variance. The switch to Salpeter does cause a dramatic spectral shift, but this is because it is extremely bottom-heavy (a large fraction of mass in low mass stars), meaning it produces much less light for a given unit of star formation mass. This is consistent with the spectra comparison in Figure 9.
@@
Assuming we using either a Chabrier or Kroupa IMF, extremely young ($10^{7}$ Gyr) and old ($10^{10}$ Gyr) stellar population spectra uncertainties are mildly dominated by stellar spectrum choice (but isochrones are not far behind). The scatter between these is minimised at a few hundred Myr, and slowly rises beyond that to a typical 5–7 % variation. The least disagreement is seen between MILES and BaSeL stellar spectra, likely because both of these are empirically constructed. C3K agrees more with these two empirical libraries than PHOENIX (Husser). Given the excellent parameter coverage of C3K (intrinsically much larger and well sampled than MILES or BaSeL), this result partly informs our decision to recommend C3K as our fiducial ‘base’ stellar spectrum library. As with the isochrones above, SSPs using all of the discussed stellar spectra have been pre-generated with ProGeny. Anybody carrying out a detailed study of SEDs with ProSpect should consider using the other stellar spectrum combinations too.
An obvious issue from an inference perspective is we do not know the truth re isochrones, stellar spectra, and/or IMF, and it is quite possible that a combination of the ‘wrong’ isochrone and ‘wrong’ IMF combine in a manner that provides a better overall fit to our data. The above findings suggest that much better fidelity in isochrones and atmospheric modelling would be needed before strong statements about IMF behaviour could be conclusively demonstrated in any absolute sense. Within a consistent SED software environment (e.g. ProSpect) relative statements about the impact of the IMF can be made however. The impact of using different SSPs (including our new ProGeny ones) is explored in depth in our Bellstedt & Robotham (in prep.) companion paper.
@@
Figure 12: Stellar mass remaining comparison of different isochrone and IMF choices with ProGeny(using Z). Other versions of this Figure (for different metallicities) can be viewed at rpubs.com/asgr.
@@
Figure 13: Ratio of stellar mass remaining for different isochrone and IMF choices (using Z). Ch = Chabrier, Kr = Kroupa, Sp = Salpeter.
4.3 Literature Comparisons
@@
Table 4: Basic information regarding the literature SSPs we compare against. Note for the stellar spectra only the dominant library is listed, nearly all SSPs use a complex mixture of spectra to achieve adequate temperature and surface gravity parameter coverage. Note M05 only has a Kroupa IMF variant.
@@
BPASS STARS C3K/BaSeL Chabrier 6 – 11 51 -3.3 – 0.3 13 Stanway & Eldridge (2018)
@@
The most fundamental low-level comparison we can make between SSPs is by analysing the actual spectra produced for specific combinations of metallicity and stellar evolutionary age. Figue 14 presents such spectra for solar metallicities, covering 100,000 yr through to 10 Gyr. For reasons not entirely clear, FSPS when run with MIST isochrones produces very unusual spectra for ages below $10^{7}$ years, with a much flatter spectral shape than seen for any other SSP. These unusual features are not present when running in the same manner but specifying Padova isochrones with FSPS (using the original Fortran software), so it would seem to be a specific issue with the younger MIST isochrones. For that reason we prefer to present FSPS using Padova isochrones for the following comparisons.
Figure 14: Comparison of Z spectra at different stellar population ages for different SSPs. The ages presented are log spaced with 1 dex intervals, covering 100,000 yr through to 10 Gyr. Zoomed in versions of the Figure (covering 1,000 – 10,000 Å), and alternative metallicity versions of this Figure can be viewed in Appendix C. The vertical dashed lines indicate the $u$ and Ks band central wavelengths, common limits for SED analysis work. The dominant isochrone used is listed in square brackets, where Mi = MIST, Pd = Padova, PC = PARSEC, St = STARS, M = Maraston, Ba = BaSTI.
@@
A consistent observation is that predictions for SSP spectra below 1,000 Å are highly variable between specific SSPs for age much above a million years, with sometimes many dex of disagreement re the detail of spectra in this regime. The reason for this is complex, but the general issue is the differing treatment of hot remnant stars, horizontal branch (HB) stars, planetary nebulae (PN), and helium burning phases by different isochrones and/or SPLs. This is clear when observing the M05 SSP that has no remnant stars and different treatment of HB stars to standard isochrones: we see a sharp truncation of UV light for all SSPs $10^{8}$ years and older with no evidence of any UV upturn at all. It is relatively rare that these parts of the SED are interrogated with lower redshift surveys since even UV facilities like GALEX only probe down to 1350 Å in the FUV. However, by redshift 0.3 the rest frame GALEX FUV filter starts to cover the 1,000 Å feature seen here, and much beyond this even typical optical filters will be measuring flux in this region.
@@
Redder than this 1000 Å feature the SSPs are remarkably consistent over 5 dex in age, with the M05 SSPs being the most obvious outlier below the 4000 Å break. Beyond that observation, the next most significant variation is seen in the near infrared (1 micron and redder) where the various AGB phases become important for ages older than $\sim 10^{7}$ years (especially above $10^{8}$ years) and have complex modelling differences at the isochrones level (as noted in Section 3.1) and decisions regarding the best (or perhaps more accurately, least bad) method for matching these stellar evolution phases to the most appropriate stellar spectra. Given the huge wealth of data, it is not surprising that models collectively agree best in the optical regime for giga-year and older stellar populations. This represents the dominant spectral feature in a huge number of passive galaxies and globular clusters, and is a common target when checking the validity of a given SSP.
@@
An additional difference we note between ProGeny (with MIST isochrones) and other SSPs can be seen in the lowest age (100,000 yr) top-left panel. Red-wards of about 1 micron there is a significant (almost 1 dex) excess to the classic black-body power-law slope you might expect (and as is seen for other SSPs). This feature is entirely cause by the tracking of pre-main sequence stars in MIST, which produces an additional cool ($\sim 2,000$ K) spectrum representing the thermal emission from collapsing molecular clouds and proto-stars. This cool spectrum is added onto the dominate spectrum of extremely hot O and B stars on the main sequence which have typical temperatures $\sim 50,000$ K. While it looks dramatic when plotted like this, the actual flux contribution in a typical galaxy SED is very small, and this difference in behaviour has little impact on composite (or complex) stellar population spectra in general.
@@
Figure 15: Stellar mass remaining comparison of different isochrone and IMF choices with ProGenyversus a compendium of literature SSPs (using Z). The dominant isochrone used is listed in square brackets, where Mi = MIST, Pd = Padova, PC = PARSEC, St = STARS, M = Maraston, Ba = BaSTI
@@ 5 Conclusions and Future Outlook
Here we have, in brief, introduced the new R based ProGeny SPL package that allows for the rapid and flexible generation of SSPs for use in particular with ProSpect, but potentially any other SED generating or fitting software. The package is thoroughly documented, with relevant examples for all user-level functions, and longer form vignettes demonstrating sophisticated types of analysis and comparison111111rpubs.com/asgr.
@@
Basic comparisons show that ProGeny is spectrally consistent with other popular SPLs when using similar isochrones and stellar spectra, with understandable variations in a few key domains (e.g. UV upturn) due to our ProGeny isochrones incorporating different physical processes to some published SPLs, even nominally identical ones. This is not surprising since isochrones with the same name (e.g. MIST, PARSEC, BaSTI etc) receive regular updates to their physics over time, and the literature is lacking in clear version control for SPLs and/or SSPs. The largest variations in spectral output are noted when switching isochrones ($\sim 60$%), followed by the stellar spectra ($\sim 20$%) and finally IMFs ($\sim 10$%, between Chabrier and Kroupa variants). This is consistent with previous literature on the topic.
@@
To create a fiducial set of stellar spectra using the recommended combination of MIST for the isochrones, C3K (for the base stellar spectra) and a Chabrier IMF takes 5s at low resolution (matching BC03) and 18s at high resolution (matching CB19) using a modern Apple MacBook. These speeds make it realistic to create SSPs on-the-fly, and for many use cases (since the fitting time will dominate the processing time with ProSpect) there is no need to pre-generate static SSPs. This is particularly useful when exploring variations of the input ingredients (like the IMF) since exhaustively generating all potential combinations is onerous in terms of disk space and book-keeping.
@@
As part of this work, we introduce a new standardised FITS based format for defining all the characteristics of stellar spectra to be incorporated into a new SPL/SSP. We noted while constructing the default input stellar spectra that every single library used its own format, ranging from pure binaries, ASCII text files, and FITS files of many different formats. While it is true we have added ‘yet another standard’, the one we propose is a minimal combination of the best characteristics of the data we processed in a highly open and well used binary (FITS) format.
@@
A companion paper (Bellstedt & Robotham in prep.) investigates the impact of varying the available ProGeny isochrones, stellar spectra, and IMFs when using the SSPs with ProSpect to fit GAMA galaxies. Longer term the aim is to establish the most plausible fiducial model, i.e. what combination of isochrones, stellar spectra and IMF (including evolving varieties) best reproduce a self consistent model of galaxy evolution in the Universe. Key to this will be investigating the implied cosmic star formation history and metallicity evolution implied by fitting galaxies at low (GAMA) and moderate (WAVES) redshift, and comparing this to ‘in-situ’ measures at earlier epochs (heavily inspired by Bellstedt et al., 2020).
@@
As with any SPL, ProGeny has certain limitations in its current form. As discussed in the main paper, we do not have explicit treatment for certain evolutionary phases that are not always well captured by traditional isochrones. The most notable features lacking being detailed modelling of the horizontal branch phases (where spectral outputs from both the core and the outer envelope become important), and planetary nebulae (where the PN can be considered as a very extended photosphere). ProGeny is not dramatically limited despite these exclusions, producing SSPs that are largely consistent with SPLs that do have such additional handling (most notably Bruzual & Charlot, 2003; Plat et al., 2019). In any case, these limitations could be overcome by ProGeny tracking multiple stellar components when relevant. We note that MESA/MIST does explicitly track the properties of the core also, the limitation then being the adoption of appropriate spectra.
Near-term planned enhancements for ProGeny include the separate treatment of remnant stars in order to make better use of isochrones that do not track these phases in detail. Until these additional tracks have been incorporated, general purpose SED modelling should probably use the MIST isochrones only. We also plan to incorporate a wider variety of isochrone libraries, where the impact of stellar rotation, $\alpha$/Fe enhancement and potentially binary evolution (with BPASS; Stanway & Eldridge, 2018) can be explored in systematic detail. To self-consistently account for more exotic forms of stellar rotation and $\alpha$/Fe enhancement the appropriate stellar spectra would also need to be added, which means such extensions are likely to happen slowly (multiple years). The ultimate hope is many of these extensions will come from the wider community, with all aspects of ProGeny being available under a permissive L-GPL3 license.
Data Availability
@@
  • Long-form examples (vignettes) that generate many of the Figures presented in this work available at ASGR’s RPubs121212rpubs.com/asgr.
  • Script to generate all 45 SSPs used in this work, also at ASGR’s RPubs.
  • Seven pre-generated ProSpect format (FITS) SSPs that use MIST isochrones, C3K base stellar spectra, and the following IMFs:
@@
    • Metallicity-Evolving Kroupa (0.1 – 100 M; default ProGeny evolution arguments)
@@
    • Metallicity-Evolving Lacey (0.1 – 100 M; default ProGeny arguments)